### 08 Jan 2024 v1

@@ Abstract
We propose GrainGNN, a surrogate model for polycrystalline grain formation under rapid solidification conditions in metal additive manufacturing. Such grain formation problems are modeled by a multicomponent partial differential equation PDE with moving interfaces. The inherent randomness of the PDE initial conditions (grain seeds) necessitates ensemble simulations to predict microstructure statistics, e.g., grain size, aspect ratio, and crystallographic orientation. Currently such ensemble simulations are prohibitively expensive and surrogates are necessary.
@@
We evaluate GrainGNN by comparing it to high-fidelity phase field simulations for in-distribution and out-of-distribution grain configurations for solidification under laser power bed fusion conditions. GrainGNN results in 80%–90% pointwise accuracy; and nearly identical distributions of scalar quantities of interest (QoI) between phase field and GrainGNN simulations compared using Kolmogorov-Smirnov test. GrainGNN’s inference speedup (PyTorch on single x86 node) over a high-fidelity phase-field simulation (CUDA on a single NVIDIA A100 GPU) is 300$\times$–2000$\times$ for 100-initial grain problem. Further, using GrainGNN, we model the formation of 11,600 grains in 220 seconds on a single CPU core.
@@ 1 Introduction
Let $\boldsymbol{\theta}(x,y,z,t)$ be the crystal orientation vector as function of space $(x,y,z)$ and time $t$. We assume we’re given high fidelity model $\mathscr{F}$ such that $\boldsymbol{\theta}(x,y,z,t)=\mathscr{F}(\boldsymbol{p},\boldsymbol{\xi},L)$. Here $\boldsymbol{p}$ encodes the meltpool temperature field using two scalars, $G$ is scalar metric of the temperature gradient and $R$ a normalized cooling rate (pooling velocity); $\boldsymbol{\xi}$ parameterizes the initial microstructure $\boldsymbol{\theta}(x,y,z=0,t=0)$ at the solid-liquid-interface (SLI); and $L=(L_{x},L_{y},L_{z})$ paramaterizes the meltpool size, which we idealize as a 3D rectangular volume with grain epitaxial growth aligned with the $z$ axis. In this context, $\boldsymbol{\theta}$ is a piecewise constant function and grains are defined as connected regions with the same $\boldsymbol{\theta}$; $\boldsymbol{\xi}$ is a random field, and quantities of interest (QoIs) that depend on $\boldsymbol{\theta}$ can only be statistically characterized and require expectations over $\boldsymbol{\xi}$.
@@
Evaluating $\mathscr{F}$ can be extremely expensive as it involves multicomponent, multiscale PDEs with moving interfaces. Furthermore, 3D printing involves a large a amount of meltpool simulations: a typical part is cm-long in the meltpool dimensions are $\mathcal{O}(100\mu$m)))). In this context, we seek to find a surrogate model $\widetilde{\boldsymbol{\theta}}=\widetilde{\mathscr{F}}(\boldsymbol{p},\boldsymbol{\xi},L)$ that approximates $\boldsymbol{\theta}$ and the statistics of the quantities of interest. We also require that $-~\widetilde{\mathscr{F}}$ is much cheaper to compute than $\mathscr{F}$.
@@
Background and significance: High fidelity models are important tools for predicting grain structure formation in rapid solidification conditions for metal additive manufacturing (AM) [1]. Common numerical methods include phase field models [2, 3, 4, 5, 6], cellular automata [7, 8, 9], and kinetic Monte Carlo methods[10] . However, these methods are computationally expensive as they require fine spatial and temporal discretizations. Furthermore, due to the stochastic nature of AM process, ensemble simulations  [11, 12, 13, 6] are required to statistically characterize quantities of interest (QoIs) like grain size, aspect ratio, and misorientation [14, 15, 16]. Tasks such as process control and optimization require repeated invocation of high-fidelity models under variable heating source conditions, feedstock composition, and component geometry. These high computational costs have restricted the adoption of high fidelity simulations for predicting or controlling the grain structure during AM processes.
@@
Summary or proposed methodology and contributions: Here, following our work in 2D surrogates [17], we focus on epitaxial grain formation and growth ignoring dendrite-level microstructure, concentration gradients, and nucleation. The key phenomena that need to be captured at this modeling level are interface evolution and grain coarsening. The latter can be decomposed into grain face elimination and grain elimination [18, 19]. The dynamics of $\mathscr{F}$ are quite challenging to capture due the evolving interface and the topological changes due to coarsening.
@@
In GrainGNN (Section 2.2), first we reduce the representation of $\boldsymbol{\theta}(x,y,z,t)$ using aggressive coarsening in space and time. (i) We introduce a reduced graph representation or $\boldsymbol{\theta}$ that captures grain boundaries, size, edge length and other geometric features per grain (Section 2.3). (ii) Following [18], we introduce a categorization of topological events that need to be tracked (Section 2.4). (iii) We use two neural networks to model their dynamics. a regressor network to capture the interface motion; and a classification network to estimate probabilities of topology-changing events (Section 2.5). (iv) We introduce graph algorithms that combine the network output to evolve the microstructure graph topology (Section 2.6). (v) We introduce an algorithm that reconstructs $\widetilde{\boldsymbol{\theta}}(x,y,z,t)$ from the reduced graph/feature-based representation of the microstructure (Section 2.7).
@@
Ground truth data is generated using an established phase-field based model of epitaxial grain growth (Section 2.1). We remark that GrainGNN is independent of the underlying solver/formulation. We train GrainGNN assuming a probability density distribution for $\boldsymbol{\xi}$, a fixed $L=L^{0}$, and a grid-sampled $\boldsymbol{p}$ (Section 3.1). Then we evaluate the in-distribution generalization of GrainGNN in which we keep $L=L^{0}$, sample training-unseen values of $\boldsymbol{\xi}$ from the training distribution, and unseen values of $\boldsymbol{p}$ and measure pointwise errors of $\boldsymbol{\theta}-\boldsymbol{\widetilde{\theta}}$ as well errors in the statistics of QoIs (Section 3.2). We also test the out-of-distribution generalization GrainGNN with $L\neq L^{0}$ and $\boldsymbol{\xi}$ sampled from a different distribution (Section 3.3). We discuss the results, extensions, and limitations of GrainGNN in Section 4.
@@
Related work: Surrogate models provide an alternative to high fidelity simulations, with the potential for reduced computational cost. Trained on either high fidelity simulation results or experimental results, such surrogates have been successful in addressing challenges in uncertainty quantification and optimization for a range of applications. [20, 21, 22, 23]. Specifically in the area of material microstructure prediction, machine learning surrogates have received increasing attention in the past few years. Convolutional neural networks have been used for 2D and 3D microstructure image reconstruction [24, 25]. Generative adversarial networks [26] are capable of learning microstructure statistics, and can approximate structure-to-material properties forward and inverse maps [27, 28]. Recurrent neural networks (RNNs) or its subclass long short-term memory (LSTM) networks [29] have been used to rapidly predict the time evolution of the microstructure. Existing works consider a representative volume element and use LSTM networks to capture evolution statistics for use cases such as: spinoidal decomposition of binary mixtures [30, 31], brittle fracture [32], single dendrite growth [32], and grain formation [17]. Ref. [33] applied the convolutional autoencoder to map the microstructure of a two phase mixture to a latent space and used DeepONet [34] to learn the evolution.
@@
Our previous work [17] introduced GrainNN, a transformer-based LSTM to predict 2D epitaxial grain growth during solidification. GrainNN reduces the computational cost by tracking only the grains instead of each grid point. It evolves manually crafted grain shape descriptors defined on each grain. GrainNN can predict microstructure with low pointwise error while achieving significant speed up over high fidelity simulations. It also utilizes domain decomposition and rectangular-to-curvilinear domain mappings to handle systems of many grains and circular geometry. However, this 2D setting the grain coupling is only one-dimensional and perpendicular to the temperature gradient, which makes GrainNN hard to model the substantially more complex grain-grain interactions in 3D.
@@
To generalize GrainNN to 3D, we utilize graph representations of the grain structure that consist of both grains and the vertices of junctions between grains. Vertex representations have previously been used in models [35, 36, 37, 19] to track motion of the grain junction points in polycrystalline microstructures. In these models, the motion of vertices during grain growth is derived from the minimization of the isotropic or anisotropic grain boundary energy. Although for vertex models the update of a grain network is efficient and schemes for handling topological transitions are proposed [35] for grain coarsening, vertex models cannot currently describe grain evolution driven by a temperature gradient during AM solidification. A grain-centric graph structure has been to predict grain microstructure evolution during solidification [38]. This physics-embedded graph network (PEGN) approach combines classic phase-field (PF) theory into a graph representation of the grains to accelerate PF simulations. Evolution is dictated by the minimization of a PF-derived free energy functional. PEGN is able to capture grain statistics in various AM setups but loses the accuracy of PF in terms of predicting actual grain shapes, due to the lack of tracking grain boundaries and topological changes.
@@
Meanwhile, graph representations of grains have been used for graph neural networks (GNNs) [39, 40] to create static microstructure-to-material-property maps. Graph convolutions on the grain networks are used to capture the spatial variations of microstructure properties and the graph sparsity enables faster evaluations of material behavior than FOMS. Beyond applications to material microstructures, GNNs have been successful in a wide range of other applications where the data is amenable to a graph structure [41, 42, 43, 44, 45, 46, 47].
@@ 2 Methods
Fig. 1: A phase field simulation of epitaxial growth of 900 grains. Temperature gradient $G_{z}=10\ K/\mu m$ and pulling velocity $R_{z}=2\ m/s$ (a) Initial substrate with width $120\mu m$ and height $2\mu m$. $\boldsymbol{\theta}_{z}$ is the angle between the crystal orientation and the $z$-axis. (b) Phase field microstructure after 10K time steps. (c) The corresponding grain index field of (b). Each grain is assigned a unique index from 1 to 900. (d) Initial grain orientation distribution. (e) Probability distributions of grain size $d$ at $t=0$ and $t=t_{H}$. (f) Time evolution of the quantities of interest. $z$ represents the height of the current solid-liquid interface. As the height of the interface increases, the number of eliminated grains increases and the volume-weight misorientation $\Delta\boldsymbol{\theta}$ decreases.
@@
In this section we discuss the overall methodology. We start with the grain growth model, then we discuss the reduced parameterization of $\boldsymbol{\theta}$, the $\boldsymbol{\theta}$-to-graph maps, the graph dynamics, the LSTM neural networks, the graph update algorithm, and the graph-to-$\boldsymbol{\theta}$ map. We conclude with details about the phase-field solver.
2.1 Problem formulation and high fidelity discretization
@@
Here we describe the model we used to generate the training data. We use a phase-field simulation. We would like to emphasize that GrainGNN does not dependent on any particular model or numerical method for generating the grain microstructure. For example level sets, different grain formation models, or cellular automata could be used. The main assumption in our surrogate is does not capture nucleation.
@@
Regarding the particular grain formation model, we adopt a grain-scale phase-field model described in Ref. [52] ignoring nucleation.  [^1] We remark that the grain structure can be modeled with even more accurate and even more expensive dendrite resolving models [53]. Each phase field component $\phi_{\alpha}$ is associated with one of $n_{g}$ different crystalline orientations, where $n_{g}$ is the number of grains. The dynamics of $\phi_{\alpha}$ is governed by the following equation.
@@
$\tau_{\alpha}\frac{\partial\phi_{\alpha}}{\partial t}=\nabla\cdot(W_{\alpha}^{2}\nabla\phi_{\alpha})+\sum_{j=x,y,z}\partial_{j}\left[|\nabla\phi_{\alpha}|^{2}W_{\alpha}\frac{\partial W_{\alpha}}{\partial(\partial_{j}\phi_{\alpha})}\right]+\phi_{\alpha}-\phi_{\alpha}^{3}-\lambda(1-\phi_{\alpha}^{2})^{2}\frac{T-T_{M}}{L/C_{p}}$
@@
$\tau_{\alpha}$ $=\tau_{0}\left(1+3\epsilon_{k}-4\epsilon_{k}({\nabla}_{x}\phi_{\alpha}^{4}+{\nabla}_{y}\phi_{\alpha}^{4}+{\nabla}_{z}\phi_{\alpha}^{4})/|{\nabla}\phi_{\alpha}|^{4}\right),$ (2.2)
$T$ $=T_{M}+G_{z}(z-R_{z}t).$
@@
Here $L_{x}$, $L_{y}$, $L^{\prime}_{z}$ are the domain dimensions for simulation and $t_{H}$ is the time horizon; $\tau_{0}$ is the interface attachment time scale; $W_{\alpha}$ is the anisotropic interface width with magnitude $W_{0}$; $\lambda$ is the thermal coupling constant; $L$ is the latent heat; $C_{p}$ is the heat capacity; $T_{M}$ is the melting temperature; $\omega$ is a scalar interaction parameter that sets the repulsive strength between adjacent grains of different orientation [3]. $\epsilon_{k}$ is the kinetic anisotropy coefficient which is assumed to have a four-fold symmetry. We assume a given temperature field parameterized by constants $G_{z}$ and $R_{z}$ in the $z$ direction.inline]removed ${}_{z}$ subscript but maybe forgot some; check; also “in the $z$ direction” is confusing since G, R are scalars. Either you need to define G and R properly, or just treat them as constants and cite papers. We specify the specific values for these parameters in the appendix. Notice that Eq. 2.2 neglects the concentration field and only considers the evolution of phase fields with prescribed temperature fields in the rapid solidification regime. A more accurate phase field model that is fully coupled with concentration and temperature field may be needed for alloys with significant solute partitioning to the liquid and under general AM conditions.
inline]Add figure that explains initial conditions and randomness and shows examples of different conditions; discuss probability density for grain substrate inline ]Clearly explain that given an alloy composition, the main free parameters that vary during the AM process are $G$, $R_{z}$; define the term “process parameters”. And state that we want to learn a surrogate model that predicts quantities of interest for different values of initial conditions. Say that all other parameters are alloy dependent. But the question is does non-dimensionalization reduce the number of these parameters?

### 31 Jan 2024 v2

@@ Abstract
We propose GrainGNN, a surrogate model for the evolution of polycrystalline grain structure under rapid solidification conditions in metal additive manufacturing. High fidelity simulations of solidification microstructures are typically performed using multicomponent partial differential equations (PDEs) with moving interfaces. The inherent randomness of the PDE initial conditions (grain seeds) necessitates ensemble simulations to predict microstructure statistics, e.g., grain size, aspect ratio, and crystallographic orientation. Currently such ensemble simulations are prohibitively expensive and surrogates are necessary.
@@
We evaluate GrainGNN by comparing it to high-fidelity phase field simulations for in-distribution and out-of-distribution grain configurations for solidification under laser power bed fusion conditions. GrainGNN results in 80%–90% pointwise accuracy; and nearly identical distributions of scalar quantities of interest (QoI) between phase field and GrainGNN simulations compared using Kolmogorov-Smirnov test. GrainGNN’s inference speedup (PyTorch on single x86 CPU) over a high-fidelity phase field simulation (CUDA on a single NVIDIA A100 GPU) is 150$\times$–2000$\times$ for 100-initial grain problem. Further, using GrainGNN, we model the formation of 11,600 grains in 220 seconds on a single CPU core.
@@ 1 Introduction
Fig. 1: Time evolution of the solidified microstructure of stainless steel 316L predicted by a phase field model and GrainGNN. (a) Phase field simulation. (b) GrainGNN predictions. $z_{l}$ is the current height of the solid-liquid interface (SLI). The domain size is $(120\mu m,120\mu m,50\mu m)$. The microstructure of 900 grains at $z_{l}=2\mu m$ is the input to both phase field simulation and GrainGNN. We apply a temperature field with $G=10K/\mu m$ and $R=2m/s$. The grains are growing in the $z$-direction and stopped at $z_{l}=50\mu m$. The elapsed simulation time for the phase field model is 48 minutes on one Nvidia A100 GPU. The GrainGNN inference time is 19 seconds on one Intel x86 6-core CPU.
@@
Let $\boldsymbol{\theta}(x,y,z,t)$ be the crystal orientation vector as a function of space $(x,y,z)$ and time $t$. We assume we are given a high fidelity model $\mathscr{F}$ such that $\boldsymbol{\theta}(x,y,z,t)=\mathscr{F}(\boldsymbol{p},\boldsymbol{\xi},L)$. Here $\boldsymbol{p}$ encodes the meltpool temperature field using two scalars, $G$ and $R$, where $G$ is a scalar metric of the temperature gradient and $R$ is a normalized cooling rate (pooling velocity); $\boldsymbol{\xi}$ parameterizes the initial microstructure $\boldsymbol{\theta}(x,y,z,t=0)$ at the solid-liquid interface (SLI); and $L=(L_{x},L_{y},L_{z})$ parameterizes the meltpool size, which we idealize as a 3D rectangular volume with grain epitaxial growth aligned with the $z$ axis (see Fig. 1). In this context, $\boldsymbol{\theta}$ is a piecewise constant function and grains are defined as connected regions with the same $\boldsymbol{\theta}$; $\boldsymbol{\xi}$ is a random field, and quantities of interest (QoIs) that depend on $\boldsymbol{\theta}$ can only be statistically characterized and require expectations over $\boldsymbol{\xi}$.
@@
Evaluating $\mathscr{F}$ can be extremely expensive as it involves multicomponent, multiscale PDEs with moving interfaces. Furthermore, additive manufacturing (AM) involves a large number of meltpool simulations: the length scale of an AM component can be of the order of meters while the meltpool dimensions are $\mathcal{O}(100\mu$m)))). In this context, we seek to find a surrogate model $\widetilde{\boldsymbol{\theta}}=\widetilde{\mathscr{F}}(\boldsymbol{p},\boldsymbol{\xi},L)$ that approximates $\boldsymbol{\theta}$ and the statistics of the quantities of interest. We also require that $-~\widetilde{\mathscr{F}}$ is much cheaper to compute than $\mathscr{F}$.
@@
Background and significance: High fidelity models are important tools for predicting grain structure formation under rapid solidification conditions existing during AM [1]. Common numerical methods include phase field models [2, 3, 4, 5, 6], cellular automata [7, 8, 9], and kinetic Monte Carlo methods[10] . However, these methods are computationally expensive as they require fine spatial and temporal discretizations. Furthermore, due to the stochastic nature of AM process, ensemble simulations  [11, 12, 13, 6] are required to statistically characterize quantities of interest (QoIs) like grain size, aspect ratio, and misorientation [14, 15, 16]. Tasks such as process control and optimization require repeated invocation of high fidelity models under variable heating source conditions, feedstock composition, and component geometry. These high computational costs have restricted the adoption of high fidelity simulations for predicting or controlling the grain structure during AM processes.
@@
Summary of proposed methodology and contributions: Here, following our work in 2D surrogates [17], we focus on epitaxial grain growth from the substrate and ignoring grain nucleation and the evolution of the solute concentration field that contributes to intragranular features of the solidification microstructure including primary and secondary arms, microsegregation, etc. The key phenomena that need to be captured at this modeling level are grain envelope evolution and grain coarsening through competitive growth. The latter can be decomposed into grain face elimination and grain elimination [18, 19]. The dynamics of $\mathscr{F}$ are quite challenging to capture due to the evolving interface and the topological changes due to coarsening.
@@
In GrainGNN (Section 2.2), first we reduce the representation of $\boldsymbol{\theta}(x,y,z,t)$ using aggressive coarsening in space and time. (i) We introduce a reduced graph representation of $\boldsymbol{\theta}$ that captures grain boundaries, size, edge length, and other geometric features per grain (Section 2.3). (ii) Following [18], we introduce a categorization of topological events that need to be tracked (Section 2.4). (iii) We use two neural networks to model their dynamics: a regression network to capture the interface motion; and a classification network to estimate probabilities of topology-changing events (Section 2.5). (iv) We introduce graph algorithms that combine the network outputs to evolve the microstructure graph topology (Section 2.6). (v) We introduce an algorithm that reconstructs $\widetilde{\boldsymbol{\theta}}(x,y,z,t)$ from the reduced graph/feature-based representation of the microstructure (Section 2.7).
@@
Ground truth data is generated using an established phase-field-based model of epitaxial grain growth (Section 2.1). We remark that GrainGNN is independent of the underlying solver/formulation. We train GrainGNN assuming a probability density distribution for $\boldsymbol{\xi}$, a fixed $L=L^{0}$, and a grid-sampled $\boldsymbol{p}$ (Section 3.1). Then we evaluate the in-distribution generalization of GrainGNN in which we keep $L=L^{0}$, sample training-unseen values of $\boldsymbol{\xi}$ from the training distribution, and unseen values of $\boldsymbol{p}$ and measure pointwise errors of $\boldsymbol{\theta}-\boldsymbol{\widetilde{\theta}}$ as well errors in the statistics of QoIs (Section 3.2). We also test the out-of-distribution generalization GrainGNN with $L\neq L^{0}$ and $\boldsymbol{\xi}$ sampled from a different distribution (Section 3.3). We discuss the results, extensions, and limitations of GrainGNN in Section 4.
@@
Related work: Surrogate models provide an alternative to high fidelity simulations, with the potential for reduced computational cost. Trained on either high fidelity simulation results or experimental results, such surrogates have been successful in addressing challenges in uncertainty quantification and optimization for a range of applications [20, 21, 22, 23]. Specifically in the area of material microstructure prediction, machine learning surrogates have received increasing attention in the past few years. Convolutional neural networks have been used for 2D and 3D microstructure image reconstruction [24, 25]. Generative adversarial networks [26] are capable of learning microstructure statistics, and can approximate structure-to-material properties forward and inverse maps [27, 28]. Recurrent neural networks (RNNs) or its subclass long short-term memory (LSTM) networks [29] have been used to rapidly predict the time evolution of the microstructure. Existing works consider a representative volume element and use LSTM networks to capture evolution statistics for use cases such as: spinoidal decomposition of binary mixtures [30, 31], brittle fracture [32], single dendrite growth [32], and grain formation [17]. Ref. [33] applied the convolutional autoencoder to map the microstructure of a two phase mixture to a latent space and used DeepONet [34] to learn the evolution.
@@
Our previous work [17] introduced GrainNN, a transformer-based LSTM to predict 2D epitaxial grain growth during solidification. GrainNN reduces the computational cost by tracking only the grain boundaries instead of each grid point inside the grains. It evolves manually crafted grain shape descriptors defined on each grain. GrainNN can predict microstructure with low pointwise error while achieving significant speed up over high fidelity simulations. It also utilizes domain decomposition and rectangular-to-curvilinear domain mappings to handle systems of many grains and circular geometry. However, in this 2D setting the grain coupling is only one-dimensional and perpendicular to the temperature gradient, which makes GrainNN hard to model the substantially more complex grain-grain interactions in 3D.
@@
To generalize GrainNN to 3D, we utilize graph representations of the grain structure that consist of both grains and the vertices of junctions between grains. Vertex representations have previously been used in models [35, 36, 37, 19] to track the motion of the grain junction points in polycrystalline microstructures. In these models, the motion of vertices during grain growth is derived from the minimization of the isotropic or anisotropic grain boundary energy. Although for vertex models the update of a grain network is efficient and schemes for handling topological transitions are proposed [35] for curvature-driven grain coarsening, vertex models cannot currently describe grain evolution driven by a temperature gradient during AM solidification. A grain-centric graph structure has been used to predict grain microstructure evolution during solidification [38]. This physics-embedded graph network (PEGN) approach combines classic phase-field (PF) theory into a graph representation of the grains to accelerate PF simulations. Evolution is dictated by the minimization of a PF-derived free energy functional. PEGN is able to capture grain statistics in various AM setups but loses the accuracy of PF in terms of predicting actual grain shapes, due to the lack of tracking grain boundaries and topological changes.
@@
Graph representations of grains have been used for graph neural networks (GNNs) [39, 40] to create static microstructure-to-material-property maps. Graph convolutions on the grain networks are used to capture the spatial variations of microstructure properties and the graph sparsity enables faster evaluations of material behavior than high fidelity models. Beyond applications to material microstructures, GNNs have been successful in a wide range of other applications where the data is amenable to a graph structure [41, 42, 43, 44, 45, 46, 47].
@@ 2 Methods
Fig. 2: A phase field simulation of epitaxial growth of 900 grains. Temperature gradient $G_{z}=10\ K/\mu m$ and pulling velocity $R_{z}=2\ m/s$ (a) Initial substrate with width $120\mu m$ and height $2\mu m$. $\boldsymbol{\theta}_{z}$ is the angle between the crystal orientation and the $z$-axis. (b) Phase field microstructure after 10K time steps. The time step size is 2.41 nanoseconds. The final height of the interface is 50 $\mu m$. (c) The corresponding grain index field of (b). Each grain is assigned a unique index from 1 to 900. (d) Initial grain orientation distribution. (e) Probability distributions of grain size $d$ at $t=0$ and $t=t_{H}$. (f) Time evolution of the quantities of interest. $z$ represents the height of the current solid-liquid interface. As the height of the interface increases, the number of eliminated grains increases and the volume-weight misorientation $\Delta\boldsymbol{\theta}$ decreases.
Table 1: Notation table.
Symbol Description (Units)
$T$ temperature field
$\phi_{\alpha}$ phase field
$\mathcal{I}$ grain index field
$\boldsymbol{\theta}$ grain orientation field $(^{\circ})$
$L$ domain size, $L=\{L_{x},L_{y},L_{z}\}$
$G_{z}$ temperature gradient $(K/\mu m)$
$R_{z}$ pulling velocity $(m/s)$
$\boldsymbol{p}$ process parameters, $p=\{G_{z},R_{z}\}$
$\boldsymbol{\xi}$ initial state of the substrate
$l$ layer index
$n_{l}$ number of layers
$I_{l}$ grain index image at height $z_{l}$
$G_{l}$ graph extracted from $I_{l}$
$F_{l}$ features of graph $G_{l}$
$g$ grain index
$j$ junction index
$V_{g}$ grain vertices of graph $G$
$V_{j}$ junction vertices of graph $G$
$n_{g}$ number of grains of graph $G$
$n_{j}$ number of junctions of graph $G$
$E_{jg}$ junction-grain edges of graph $G$
$E_{jj}$ junction-junction edges of graph $G$
$N_{j}$ grain neighbors of junction $j$
$N_{g}$ junction neighbors of grain $g$
@@
In this section we discuss the overall methodology. We start with the grain growth model, then we discuss the reduced parameterization of $\boldsymbol{\theta}$, the $\boldsymbol{\theta}$-to-graph map, the graph dynamics, the LSTM neural networks, the graph update algorithm, and the graph-to-$\boldsymbol{\theta}$ map. We conclude with details about the phase field solver.
2.1 Problem formulation and high fidelity model
@@
Here we describe the model we used to generate the training data. We use a phase field simulation. We would like to emphasize that GrainGNN does not depend on any particular model or numerical method for generating the grain microstructure. For example level sets, different grain formation models, or cellular automata could be used. The main assumption in our surrogate is does not capture nucleation.
@@
Regarding the particular grain formation model, we adopt a grain-scale phase field model described in Ref. [52] and ignore nucleation [^1] We remark that the model we use neglects the concentration field and only considers the evolution of phase fields with temperature fields in the rapid solidification regime. The grain structure can be modeled with even more accurate and even more expensive dendrite-resolving models [53].. Each phase field component $\phi_{\alpha}$ is associated with one of $n_{g}$ different crystalline orientations, where $n_{g}$ is the number of grains. The dynamics of $\phi_{\alpha}$ is governed by the following equation.
@@
$\tau_{\alpha}\frac{\partial\phi_{\alpha}}{\partial t}=\nabla\cdot(W_{\alpha}^{2}\nabla\phi_{\alpha})+\sum_{j=x,y,z}\partial_{j}\left[|\nabla\phi_{\alpha}|^{2}W_{\alpha}\frac{\partial W_{\alpha}}{\partial(\partial_{j}\phi_{\alpha})}\right]+\phi_{\alpha}-\phi_{\alpha}^{3}-\lambda(1-\phi_{\alpha}^{2})^{2}\frac{T-T_{M}}{L_{f}/C_{p}}$
[…]
$\phantom{BCs:}\phi_{\alpha}(y=0)=\phi_{\alpha}(y=L_{y})$ (2.1c)
$\phantom{BCs:}{\nabla}_{z}\phi_{\alpha}(z=0,L^{\prime}_{z})=0$ (2.1d)
@@
$W_{\alpha}$ $=W_{0}\left(1-3\epsilon_{c}+4\epsilon_{c}({\nabla}_{x}\phi_{\alpha}^{4}+{\nabla}_{y}\phi_{\alpha}^{4}+{\nabla}_{z}\phi_{\alpha}^{4})/|{\nabla}\phi_{\alpha}|^{4}\right),$ (2.2a)
$\tau_{\alpha}$ $=\tau_{0}\left(1+3\epsilon_{k}-4\epsilon_{k}({\nabla}_{x}\phi_{\alpha}^{4}+{\nabla}_{y}\phi_{\alpha}^{4}+{\nabla}_{z}\phi_{\alpha}^{4})/|{\nabla}\phi_{\alpha}|^{4}\right),$ (2.2b)
$T$ $=T_{M}+G_{z}(z-R_{z}t).$ (2.2c)
@@
Here $L_{x}$, $L_{y}$, $L^{\prime}_{z}$ are the domain dimensions for simulation and $t_{H}$ is the time horizon; $\tau_{0}$ is the interface attachment time scale; $W_{0}$ is the width of the anisotropic interface; $\lambda$ is the thermal coupling constant; $L_{f}$ is the latent heat; $C_{p}$ is the heat capacity; $T_{M}$ is the melting temperature; $\omega$ is a scalar interaction parameter that sets the repulsive strength between adjacent grains of different orientations [3]. $\epsilon_{c}$ and $\epsilon_{k}$ are the capillary and kinetic anisotropy coefficients respectively, which are assumed to have a four-fold symmetry. In Eq. 2.2c, we assume the temperature gradient $G_{z}$ and the pulling velocity $R_{z}$ are aligned with the $z$-axis and they are constants during the simulation. Under this temperature profile, grains are growing in the $z$-direction. We name $\boldsymbol{p}=\{G_{z},R_{z}\}$ the process parameters. $\boldsymbol{p}$ are the main free parameters during AM processes given the alloy with its material parameters. We specify the values for the material parameters in the appendix. The discretization of Eq. 2.1a is discussed in Section 2.8.
We use no-flux boundary conditions at the top and bottom surface ($z=0,z=L^{\prime}_{z}$) and periodic boundary conditions at the four sides of the domain. We initialize $\phi_{\alpha}$ as follows. Assume that the initial SLI is at $z=z_{0}$, we partition the $z=z_{0}$ plane into $n_{g}$ regions. For phase field $\alpha$, if point $(x,y)$ is in region $\Omega_{\alpha}$ we use a $\tanh$ function to create a smooth transition from solid (1) to liquid (-1) in the $z$-direction; if point $(x,y)$ is outside of $\Omega_{\alpha}$, we set $\phi_{\alpha}$ to -1. $\boldsymbol{\xi}$ parameterizes the initial realization of $\Omega_{\alpha}$. Here we use the Voronoi diagram [6] which creates the Voronoi tessellations from a set of seed points. We let $\boldsymbol{\xi}=\{\boldsymbol{x}^{0}_{j}\}^{n_{j}}_{j=1}\cup\{\boldsymbol{\theta}_{g}\}^{n_{g}}_{g=1}$, where $\boldsymbol{x}^{0}_{j}$ are the vertices of the Voronoi diagram and $\boldsymbol{\theta}_{g}$ are the grain orientation for each phase field. With a crystal orientation angle $\boldsymbol{\theta}_{g}$, the anisotropy functions Eq. 2.2a and Eq. 2.2b are modified by replacing all spatial derivatives by the derivatives with respect to the rotated coordinate system ($x^{\prime}$, $y^{\prime}$, $z^{\prime}$) with angle $\boldsymbol{\theta}_{g}$ [54]. For each phase field simulation, $\boldsymbol{\xi}$ is randomly generated. The orientation of each grain is sampled from the unit sphere. If $\boldsymbol{d}$ is the vector representing the grain’s orientation, we first sample $\boldsymbol{d}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ and set $\boldsymbol{d}=\boldsymbol{d}/|\boldsymbol{d}|_{2}$. Samplings of $\boldsymbol{x}^{0}_{j}$ are discussed in Section 3.1.
Fig. 2 shows an example of a simulation with domain size $(120\mu m,120\mu m,50\mu m)$. The interface width $W_{0}=0.1\mu m$ and the mesh size $dx=0.08\mu m$ [6]. Therefore the grid size is $1500\times 1500\times 625$. The initial conditions for grain microstructure in the substrate are generated by sampling $\boldsymbol{x}^{0}_{j}$ from a uniform distribution and using a Voronoi diagram with periodic boundary conditions. In Fig. 2a, the substrate has 900 grains. In this simulation $G_{z}$ and $R_{z}$ are $10K/\mu m$ and $2m/s$, respectively. The time step size is 2.42 nanoseconds. Fig. 2b is the resulting phase field microstructure after 10K time steps, which required approximately 48 minutes on one Nvidia A100 GPU.